# Learning-based Multi-View Stereo: A Survey

### v2

@@ 2.6 Evaluation Metrics
2D Metrics: 2D depth-based metrics are commonly used by online MVS methods [25, 30] to evaluate the depth maps [68, 69]. The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$). For more details, please refer to supplementary.
3D Metrics: 3D point cloud evaluation is widely used by offline MVS methods [15, 36]. Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated. The metrics are Precision/Accuracy, Recall/Completeness and F-score. For more details, please refer to supplementary.
3 Supervised Method with Depth Estimation
This section mainly introduces supervised learning-based MVS methods with depth estimation. A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec. 3.1), cost volume construction (Sec. 3.2), cost volume regularization (Sec. 3.3) and depth estimation (Sec. 3.5). The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig. 3 and Fig. 4, which are the representative methods of online and offline MVS methods respectively.
@@ 3.1 Feature Extraction
Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72]. For online MVS methods, feature extraction networks are usually used in line with the real-time operation goal. DeepVideoMVS [30] combines MNasNet [73], which is lightweight and has low latency, with FPN. SimpleRecon [34] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy. For offline MVS methods, MVSNet [15] uses a stacked 8-layer 2D CNN to extract deep features for all the images. Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [31, 38, 39] or multi-scale RGB images [75, 76]. Recently, many following works have been paying more attention to feature extraction to improve the representation power of deep features. [77, 78, 79] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture. CDS-MVSNet [80] uses a dynamic scale network, CDSFNet, to extract the discriminative features by analyzing the normal curvature of the image surface. GeoMVSNet [76] proposes a geometry fusion network that fuses the image features with coarse depth maps. With FPN as the main feature extractor, WT-MVSNet [81] and MVSFormer [82], further introduces Vision Transformers [83, 84, 85] for image feature enhancement. ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
@@ 3.2 Cost Volume Construction
For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3.
@@ 3.2.1 Online MVS
To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$. MVDepthNet [25] and GP-MVS [60] compute the per-pixel intensity difference between the reference and each source view as matching cost. If there is more than one source view, the cost volumes are averaged. Instead of intensity difference, Neural RGB-D [87] computes deep feature difference. DeepVideoMVS [30] and MaGNet [88] calculate the cost from pixel-wise correlation between the reference feature and the warped source feature. In addition to the dot product between reference image features and warped source image features, SimpleRecon [34] further adds metadata, e.g., ray direction and relative pose, into the 4D cost volume. Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume. DoubleTake [89] further includes depth hint, the depth map rendered from the previously predicted geometry, into the cost volume.
@@ 3.2.2 Offline MVS
Offline MVS methods focus on reconstructing high-quality dense geometry with high-resolution images. To encode more matching information and improve the quality, offline methods [15, 31] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$. Since there may exist an arbitrary number of source views and serious occlusion [23], robustly aggregating matching information from all the source views is an important step. MVSNet [15] proposes a variance-based cost metric as follows:
@@
where $\{\mathbf{V}_{i}\}_{i=0}^{N-1}$ are (warped) feature volumes, $\overline{\mathbf{V}}$ is the average feature volume. To further reduce dimension, CIDER [35] adopts group-wise correlation [90] to compute a lightweight cost volume between reference and each warped source view. Then $N-1$ cost volumes are averaged for regularization. Without considering occlusions, these methods consider all source views equally with the averaging operation.
@@ 3.3.1 Online MVS
Figure 3: Pipeline of MVDepthNet [25]. Multiple image frames are encoded in the cost volume. MVDepthNet takes the reference image and the cost volume as the input and outputs inverse depth maps of four different resolutions. Skip connections between the encoder and decoder of the same resolution are used for the per-pixel depth estimation.
Figure 4: Pipeline of MVSNet [15]. Reference and source images go through the feature extraction network, followed by the differentiable homograph warping to construct the cost volume. A 3D U-Net is used to regularize the cost volume into a probability volume. The final depth map is estimated from the probability volume and refined with the reference image features.
@@ 3.3.2 Offline MVS
Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN.
Direct 3D CNN: Similar to stereo estimation [62, 63, 64], 3D CNN is widely used for cost volume regularization in MVS [15, 96, 97, 98, 35]. As the blueprint of learning-based offline MVS methods, MVSNet [15] adopts a 3D U-Net [71] to regularize the cost volume, which aggregates context information from a large receptive field with relatively low computation cost. It is found that the 3D regularization can capture better geometry structures, perform photometric matching in 3D space, and alleviate the influence of image distortion caused by perspective transformation and occlusions [97]. However, since 3D CNN is memory and run-time consuming, many offline methods [15, 97, 98, 35] use limited depth hypotheses and estimate depth maps at low resolution.
RNN: Instead of 3D CNN, R-MVSNet [36] sequentially regularizes 2D slices of the cost volume along the depth dimension with a convolutional GRU [99], which is able to gather spatial and uni-directional context information in the depth dimension. $D^{2}$-RMVSNet [37] augments R-MVSNet [36] with a complex convolutional LSTM [95]. AA-RMVSNet [77] further introduce an intra-view feature aggregation module for feature extraction and an inter-view cost volume aggregation module to adaptively aggregate cost volumes of different views. With sequential process of 2D slices, these methods improve the scalability for high-resolution reconstruction as well as large-scale scenes and reduce memory, however, at the cost of run-time.
Coarse-to-Fine: Predicting depth in a coarse-to-fine manner [31, 38, 75, 39, 78, 94, 100] is another solution of reducing both memory consumption and running-time. A coarse depth map is first predicted and then upsampled and refined during finer stages to construct fine details. CasMVSNet [31] constructs cascade cost volumes by warping features with reduced depth ranges around the previous coarse maps. Based on Eq. 3, the homography of differential warping at stage $k+1$ is
@@
where $d^{(k)}$ is the estimated depth value at stage $k$ and $\Delta^{(k+1)}$ is the residual depth to be determined in the current stage. Following MVSNet [15], 3D CNN is used on each stage to regularize the cost volume. UCS-Net [38] estimates uncertainty from the coarse prediction to adaptively adjust search ranges in finer stages. CVP-MVSNet [75] constructs the cost volume with a proposed optimal depth resolution of half pixel to narrow depth range in finer stages. EPP-MVSNet [101] introduces an epipolar-assembling module to assemble high-resolution information into cost volume and an entropy-based process to adjust depth range. TransMVSNet [79] uses a Feature Matching Transformer for robust long-range global context aggregation within and across images. WT-MVSNet [81] uses a window-based Epipolar Transformer for enhanced patch-to-patch matching, and a window-based Cost Transformer to better aggregate global information.
@@ 3.4 Iterative Update
Some methods [39, 102] combine iterative PatchMatch [105] with deep learning. PatchMatch algorithm is widely used in many traditional MVS methods [106, 18, 33, 19]. The PatchMatch algorithm mainly consists of: random initialization, propagation of hypotheses to neighbors, and evaluation for choosing best solutions. After initialization, the approach iterates between propagation and evaluation until convergence. Traditional methods usually design fixed patterns for propagation. Recently, PatchmatchNet [39] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps. Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features. PatchMatch-RL [102], jointly estimates depth, normal and visibility with a coarse-to-fine structure. Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in training.
Recently, RAFT [107] estimates optical flow by iteratively updating a motion field with GRU. In MVS, several methods [40, 104, 103] also adopted this approach to enhance efficiency and flexibility. IterMVS [40] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state. CER-MVS [103] and Effi-MVS [104] both embed the RAFT module in a coarse-to-fine structure, which outputs a residual that is added to the previous depth. MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores. DELS-MVS [108] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space. Based on IterMVS [40], RIAV-MVS [109] iteratively refines index fields with a 3-level GRUs. IGEV-MVS [110] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of correlation volumes.
@@ 3.5.1 Online MVS
Many methods [25, 60, 30, 34] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1, i.e., $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W\times 1}$. Then the sigmoid activation $\sigma$ is applied for normalization. Together with the predefined depth range $[d_{\textrm{min}},d_{\textrm{max}}]$ (mostly manually set), the depth map can be computed. For example, DeepVideoMVS [30] estimates depth as:
@@ 3.5.2 Offline MVS
$\hat{d}(\mathbf{p})=\sum_{i=1}^{D}d_{i}\cdot\mathbf{P}(i,\mathbf{p}),$ (8)
where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $P(i,\mathbf{p})$ is the predicted depth probability. For coarse-to-fine methods [31, 38, 75], soft argmax is applied on each stage to regress the depth maps. On the other hand, some methods [35, 39] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [35].
@@
DMVSNet [112] estimates two depth maps on multi-stages and composes the final depth map by alternating between selecting the maximum and minimum predicted depth values, which generates an oscillating depth map for improved geometry.
@@ 3.6 Depth Refinement
Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy. R-MVSNet [36] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision. Point-MVSNet [97] uses PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth. Fast-MVSNet [113] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals. PatchmatchNet [39] refines the final upsampled depth map with a depth residual network [114] and reference image feature.  [40, 104, 103] use the mask upsampling module from RAFT [107] to upsample and refine the depth map to full resolution by computing the weighted sum of depth values in a window based on the mask. Based on a coarse depth map, RayMVSNet [115, 116] aggregates multi-view image features with an epipolar transformer and use a 1D implicit field to estimate the SDF of the sampled points and the location of the zero-crossing point. GeoMVSNet [76] filters the depth map by geometry enhancement in the frequency domain. EPNet [117] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.

### v3

@@ 2.6 Evaluation Metrics
2D Metrics: 2D metrics are used to evaluate depth maps [79, 80]. The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel), and inlier ratio with threshold 1.25 ($\delta<1.25$). For more details, please refer to the supplement.
3D Metrics: 3D metrics are used to evaluate the accuracy of point cloud. Note that the reconstructed point clouds should be aligned with ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated. The metrics are Precision/Accuracy, Recall/Completeness and F-score. For more details, please refer to the supplement.
3 Supervised MVS with Depth Representation
This section mainly introduces supervised MVS methods with depth representation. As shown in Fig. 3, a typical depth map-based MVS pipeline mainly consists of feature extraction (Sec. 3.1), cost volume construction (Sec. 3.2), cost volume regularization (Sec. 3.3), and depth estimation (Sec. 3.5).
@@ 3.1 Feature Extraction
Considering efficiency, most methods use simple CNN structures to extract deep features $\left\{\mathbf{F}_{i}\right\}_{i=0}^{N-1}$ from images, e.g., ResNet [81], U-Net [82] and FPN [83].
For online MVS methods, feature extraction networks are usually chosen in line with the real-time operation goal. DeepVideoMVS [31] combines MNasNet [84], which is lightweight and has low latency, with FPN. SimpleRecon [34] and DoubleTake [85] use the first two blocks of ResNet18 [81] for the cost volume and an EfficientNet-v2 [86] encoder for the monocular image prior, which maintains efficiency and yields a significant improvement in depth map accuracy.
For offline MVS methods, MVSNet [9] uses a stacked 8-layer 2D CNN to extract deep features for all images. Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [32, 38, 39] or multi-scale RGB images [87, 88]. Recently, many following works pay more attention to feature extraction to improve the representation power of deep features. [89, 90, 91] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture. TransMVSNet [91] adpots a Feature Matching Transformer for robust long-range global context aggregation within and across images. With FPN as the main feature extractor, WT-MVSNet [92] and MVSFormer++ [93], further introduce Vision Transformers [94, 95, 96] for image feature enhancement.
@@ 3.2 Cost Volume Construction
For both online and offline MVS, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3 with depth samples $\mathbf{D}\in\mathbb{R}^{H\times W\times D}$.
@@ 3.2.1 Online MVS
To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$. The cost volume stores a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d$, which we denote as $\mathbf{C}(\mathbf{p},d)$. For example, MVDepthNet [26] and GP-MVS [67] compute the per-pixel intensity difference between the reference and each source view and average the matching costs from all source views:
$\mathbf{C}(\mathbf{p},d)=\frac{1}{N-1}\sum_{i=1}^{N-1}AD(\mathbf{I}_{0}(\mathbf{p}),\mathbf{I}_{i}(\mathbf{p}_{i}(d))),$ (6)
where $AD(\cdot,\cdot)$ denotes the absolute intensity difference, $\mathbf{p}_{i}(d)$ is the reprojected pixel computed from Eq. 5. Instead of intensity difference, Neural RGB-D [97] computes the $L_{2}$ norm of the difference between the reference and warped source features as the matching cost. DeepVideoMVS [31] computes the dot product of the reference and warped source features:
$\mathbf{C}(\mathbf{p},d)=-\frac{1}{C\times(N-1)}\sum_{i=1}^{N-1}\langle\mathbf{F}_{0}(\mathbf{p}),\mathbf{F}_{i}(\mathbf{p}_{i}(d))\rangle,$ (7)
where $\mathbf{F}_{0},\mathbf{F}_{i}$ are image features, and $C$ is the number of feature channels. Based on DeepVideoMVS, SimpleRecon [34] further introduces metadata, e.g., ray direction and relative pose, into the cost volume, which is found to increase the accuracy. DoubleTake [85] further includes depth hint, the depth map rendered from the previously predicted geometry, into the cost volume.
@@ 3.2.2 Offline MVS
To encode more matching information and improve reconstruction quality, offline MVS methods [9, 32] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d$ is associated with a matching cost of dimension $C$. Since there may exist an arbitrary number of source views and serious occlusion [24], robustly aggregating matching information from all source views is an important step. MVSNet [9] proposes a variance-based cost volume as follows:
@@
where $\mathbf{V}_{0}$ is the reference feature volume, and $\{\mathbf{V}_{i}\}_{i=i}^{N-1}$ are warped feature volumes of source views, $\overline{\mathbf{V}}=1/N\sum_{i=0}^{N-1}\mathbf{V}_{i}$ is the average feature volume. Specifically, $\mathbf{V}_{0}(\mathbf{p},d)=\mathbf{F}_{0}(\mathbf{p})$, $\mathbf{V}_{i}(\mathbf{p},d)=\mathbf{F}_{i}(\mathbf{p}_{i}(d))$. To reduce computation, CIDER [35] adopts group-wise correlation [71] to compute a lightweight cost volume between reference and warped features from N1N-1 source views. Then the individual cost volumes for different source images are averaged as the final cost volume $\mathbf{C}$. Without considering occlusions, these methods consider all source views equally with the averaging operation.
@@ 3.3.1 Online MVS
Figure 3: Pipeline of depth map-based MVS, which usually consists of feature extraction, cost volume construction via plane sweep, cost volume regularization, depth estimation, and depth refinement.
@@ 3.3.2 Offline MVS
Among most offline MVS methods that use 4D cost volumes, there are three main categories of cost volume regularization: direct 3D CNN, RNN, and coarse-to-fine.
Direct 3D CNN: Similar to stereo estimation [69, 70, 71], 3D CNN is widely used for cost volume regularization in MVS [9, 103, 104, 105, 35]. While this concept has been explored for stereo matching in works like DPSNet [103], MVSNet [9] serves as the blueprint for learning-based offline MVS methods. It adopts a 3D U-Net [82]to regularize the cost volume, which aggregates context information from a large receptive field with relatively low computation cost. It is found that the 3D regularization can capture better geometry structures, perform photometric matching in 3D space, and alleviate the influence of image distortion caused by perspective transformation and occlusions [104]. However, since 3D CNN is memory and run-time consuming, many offline methods [9, 104, 105, 35] use limited depth hypotheses and estimate depth maps at low resolution.
RNN: Instead of 3D CNN, R-MVSNet [36] sequentially regularizes 2D slices of the cost volume along the depth dimension with a convolutional GRU [106], which is able to gather spatial and uni-directional context information in the depth dimension. $D^{2}$HC-RMVSNet [37] augments R-MVSNet [36] with a complex convolutional LSTM [102]. AA-RMVSNet [89] further introduce an intra-view feature aggregation module for feature extraction and an inter-view cost volume aggregation module to adaptively aggregate cost volumes of different views. With sequential process of 2D slices, these methods improve the scalability for high-resolution reconstruction as well as large-scale scenes and reduce memory, however, at the cost of run-time [107].
Coarse-to-Fine: Predicting depth in a coarse-to-fine manner [32, 38, 87, 39, 90, 101, 108] is another solution of reducing both memory consumption and running-time. A coarse depth map is first predicted and then upsampled and refined during finer stages to construct fine details. CasMVSNet [32] constructs cascade cost volumes by warping features with reduced depth ranges around the previous coarse depth maps. Based on Eq. 4, the homography of differential warping at stage $k+1$ is:
@@
where $d^{(k)}$ is the estimated depth value at stage $k$ and $\Delta^{(k+1)}$ is the residual depth to be determined in the current stage. Following MVSNet [9], 3D CNN is used on each stage to regularize the cost volume. UCS-Net [38] estimates the uncertainty from coarse prediction to adaptively adjust search ranges in the finer stages. CVP-MVSNet [87] constructs the cost volume with a proposed optimal depth resolution of half pixel to narrow depth range in finer stages. EPP-MVSNet [109] introduces an epipolar-assembling module to assemble high-resolution information into cost volume and an entropy-based process to adjust depth range. WT-MVSNet [92] uses a window-based Epipolar Transformer for enhanced patch-to-patch matching, and a window-based Cost Transformer to better aggregate global information.
@@ 3.4 Iterative Update
Some methods [39, 110] combine iterative PatchMatch [113] with deep learning. PatchMatch algorithm is widely used in many traditional MVS methods [114, 18, 30, 19]. The PatchMatch algorithm mainly consists of: random initialization, propagation of hypotheses to neighbors, and evaluation for choosing best solutions. After initialization, the approach iterates between propagation and evaluation until convergence. Traditional methods usually design fixed patterns for propagation. Recently, PatchmatchNet [39] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps. PatchMatch-RL [110], jointly estimates depth, normal and visibility with a coarse-to-fine structure. Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in training.
Recently, RAFT [115] estimates optical flow by iteratively updating a motion field with GRU. In MVS, several methods [40, 112, 111] also adopted this approach to enhance efficiency and flexibility. IterMVS [40] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state. CER-MVS [111] and Effi-MVS [112] both embed the RAFT module in a coarse-to-fine structure, which outputs a residual that is added to the previous depth. MaGNet [116] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores. DELS-MVS [117] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space. IGEV-MVS [118] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of correlation volumes. Wang et al. [119] introduce diffusion model in MVS and propose a conditional diffusion network that iteratively denoises the depth map, achieving high accuracy and efficiency.
@@ 3.5.1 Online MVS
Many methods [26, 67, 31, 34] apply encoder-decoder on the cost volume $\mathbf{C}$ and reduce the feature channel to 1 as $\mathbf{C}^{\prime}\in\mathbb{R}^{H\times W}$. Then the sigmoid activation $\sigma$ is applied for normalization. Together with the predefined depth range $d_{\textrm{min}},d_{\textrm{max}}]$ (mostly manually set), the depth map can be computed. For example, DeepVideoMVS [31] estimates depth as:
@@ 3.5.2 Offline MVS
$\hat{d}(\mathbf{p})=\sum_{i=1}^{D}d_{i}\cdot\mathbf{P}(\mathbf{p},i),$ (11)
where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $\mathbf{P}(\mathbf{p},i)$ is the predicted depth probability. For coarse-to-fine methods [32, 38, 87], soft argmax is applied on each stage to regress the depth maps. On the other hand, some methods [35, 39] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [35].
@@ 3.6 Depth Refinement
Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy. DeepMVS [121] apply DenseCRF [122] to the coarse disparity map to encourage the pixels which are spatially close and with similar colors to have closer disparity predictions. R-MVSNet [36] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision. Point-MVSNet [104] uses PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth. Fast-MVSNet [123] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals. PatchmatchNet [39] refines the final upsampled depth map with a depth residual network [124] and reference image feature.  [40, 112, 111] use the mask upsampling module from RAFT [115] to upsample and refine the depth map to full resolution by computing the weighted sum of depth values in a window based on the mask. Based on a coarse depth map, RayMVSNet [125, 126] aggregates multi-view image features with an epipolar transformer and use a 1D implicit field to estimate the SDF of the sampled points and the location of the zero-crossing point. GeoMVSNet [88] filters the depth map by geometry enhancement in the frequency domain. EPNet [127] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.